# A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation

### v1

@@ 3.2 Loss Functions
We use five loss functions to train the proposed method. Two are standard CycleGAN loss functions, i.e., adversarial loss [20] and cycle consistency loss [89]. Inspired from [89] we also use an identity loss [76]. Finally, we introduce two novel loss functions for HDR reconstruction, i.e., contrastive and semantic segmentation loss.
The loss is calculated on tone-mapped versions of the reconstructed and real HDR images. The tone-mapping is performed on the basis of the $\mu$-law [34]. This is done to avoid the high intensity pixels of HDR images that can distort the loss calculation. The tone-mapping operator $T$ can be represented as follows:
@@
Cycle Consistency Loss. The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each another. Hence, we also incorporate a cycle consistency loss to prevent mutual contradiction of the learned mappings $G_{Y}$ and $G_{X}$. For each image $x_{i}$ from the LDR domain, the cycle of reconstruction i.e., from LDR to HDR and then back to LDR must result back in the original image $x_{i}$. Hence, we can define the forward cycle consistency as: $x_{i}\rightarrow G_{Y}(x_{i})\rightarrow G_{X}(G_{Y}(x_{i}))\approx x_{i}$. Similarly, the backward cycle consistency can be represented as: $y_{j}\rightarrow G_{X}(y_{j})\rightarrow G_{Y}(G_{X}(y_{j}))\approx y_{j}$. The loss can be formulated as:
@@
where $\|\cdot\|_{1}$ is the L1 norm (cf., Supplementary for visual depiction of the cycle consistency loss).
Identity Loss. For LDR $\leftrightarrow$ HDR translation tasks, we also find that adversarial and cycle consistency formulations alone cannot preserve the color and hue information. This is due to incorrect mapping of color shades from LDR to HDR domains by the generators stemming from the underlying difference in dynamic ranges. Therefore, we force the generators to replicate an identity mapping by providing target domain images. This loss can be expressed as:
$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (5)
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [8]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(y)$ we first define the cosine similarity between them as follows:
@@
where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs. This loss is calculated in both the forward and backward cycles with the input and output being swapped. This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (cf., Supplementary for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection of Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [8]. We choose equalized images instead of original LDR because segmentation on low-light or extreme brightness images does not yield good results. We use Segment Anything (SAM) [42] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images (cf., Supplementary for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as:
$\text{IoU}_{c}=\frac{x_{c}\cap y_{c}}{x_{c}\cup y_{c}},$ (8)
where $x_{c}$ and $y_{c}$ represent the segmentations for class $c$ in image $x$ and $y$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$. The mean IoU over all segmentation classes can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (9)
We represent the semantic segmentation loss as:
$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (10)
$\mathcal{L}_{\text{full}}=\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}$ (11)
Final Loss. The full objective is expressed in Eq. (11), where $\lambda$ scales the relative importance of the cycle consistency and identity loss. $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [89]. $\alpha$ and $\beta$ are weights for the contrastive and semantic segmentation loss, respectively. Both values are set to $2$ (cf., Supplementary for ablation experiments for $\alpha$ and $\beta$).
4 Experiments
4.1 Implementation
The proposed method is implemented in PyTorch [^1] https://pytorch.org/ on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD. We used Adam optimizer [41] to train the models for $170$ epochs. Batch size of $1$ is expected to work well in cycle consistency models [89] however, since we are using contrastive loss on positive and negative pairs, the batch size is set to $4$. We used learning rate of $4\times 10^{-4}$ for the generators and $2\times 10^{-4}$ for the discriminators. The learning rate is kept constant for the first $100$ epochs and subsequently set to decay linearly to 00. For all models, we resized the input images to $512\times 512$. The HDR images used in the text are tone-mapped with Reinhard’s operator [68].
4.2 Datasets
For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [13], NTIRE [64], and HDR-Synth & HDR-Real [51] paired datasets. HDRTV has $1235$ samples for training and $117$ for testing. The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples. HDR-Synth & HDR-Real dataset consists of $20537$ samples. Other paired datasets used in our evaluations are DrTMO [19] ($1043$ samples), Kalantari [36] ($89$ samples), HDR-Eye [43] (46 samples), and LDR-HDR Pair [33] ($176$ samples). We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity. We perform $80/20$ split for training and testing unless specified otherwise. For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR. For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [8] for datasets with only single-exposed LDR images.
4.3 Metrics
We use four metrics to report the results. High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [55] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model. Structural Similarity Index Measure (SSIM) [79, 80, 81] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information. Peak Signal-to-Noise Ratio (PSNR) [26] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^2] PSNR is computed on $\mu$-law based tone-mapped HDR images.. Learned Perceptual Image Patch Similarity (LPIPS) [87] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge based contributions. This metric aligns with the perceptual judgement of humans.
Table 2: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with supervised (gray) and unsupervised/weakly-supervised/self-supervised (black) learning methods trained on the paired datasets HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51].
[…]
HDRCNN [18] 20.12 0.724 0.422 60.12
DrTMO [19] 20.43 0.711 0.412 60.81
ExpandNet [56] 21.44 0.854 0.451 58.12
FHDR [39] 20.84 0.841 0.307 62.45
SingleHDR [51] 21.01 0.851 0.402 64.21
Two-stage [1] 34.29 0.856 0.287 61.78
HDR-GAN [63] 30.11 0.912 0.411 65.01
KUNet [77] 37.21 0.974 0.051 62.14
Ghost-free [52] 40.32 0.966 0.043 63.02
ArtHDR-Net [6] 37.12 0.961 0.321 63.43
HistoHDR-Net [5] 38.21 0.968 0.301 65.15
UPHDR-GAN [49] 39.98 0.965 0.044 63.54
PSENet [62] 27.35 0.856 0.274 62.89
SingleHDR(W) [45] 30.79 0.952 0.055 62.95
SelfHDR [88] 39.51 0.972 0.037 64.77
Ours 40.01 0.979 0.021 68.41
Table 3: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with UPHDR-GAN trained on unpaired datasets.
@@ 4.3 Metrics
Ours 43.92 0.989 0.020 69.32
5 Results
5.1 HDR Reconstruction
For an extensive comparison we choose a combination of methods that utilize different learning paradigms including, HDRCNN [18], DrTMO [19], ExpandNet [56], FHDR [39], SingleHDR [51], Two-stage [1], KUNet [77], HistoHDR-Net [5], HDR-GAN [63], ArtHDR-Net [6], Ghost-free HDR [52], SelfHDR [88], UPHDR-GAN [49], SingleHDR(W) [45], and PSENet [62]. These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised/weakly-supervised/self-supervised, unpaired learning, and knowledge-conditioned.
Table 4: LDR reconstruction (tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with the state-of-the-art one-mapping operators.

### v2

@@ 3.2 Loss Functions
We introduce six loss functions to train the proposed method. We introduce three novel loss functions for HDR reconstruction - LLM-based, contrastive, and semantic segmentation loss. Two are standard loss functions, i.e., adversarial loss [25] and cycle consistency loss [117]. Inspired from [117] we also use an identity loss [93].
The calculations in all loss functions are based on tone-mapped versions of the reconstructed and real HDR images. This tone-mapping is performed on the basis of the $\mu$-law [43] and is done to avoid the high-intensity pixels of HDR images that can distort the loss calculation. The tone-mapping operator $T$ can be represented as follows:
@@
LLM Loss. Given a reconstructed HDR/LDR image, this loss is based on pixels identified by the LLM as artifacts and over/underexposed regions during Q&A sessions (using human-defined knowledge). The loss is defined as:
$\mathcal{L}_{\text{llm}}=3\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+2\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+1.5\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed. The weights on each of the three components are selected after thorough ablations as well as theoretical considerations. Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye. Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component. (Sec. S2.3 for visual depiction of the LLM loss). This loss is calculated in the backward cycle with $\hat{x}$ but only with the artifact component. The rationale behind this decision lies in the nature of the output, i.e., LDR can be overexposed or underexposed.
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [10]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as follows:
[…]
where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs. This loss is calculated in both the forward and backward cycles with the input and output being swapped. This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (Secs. S2.3 and S3.3 for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection over Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [10]. We choose equalized images instead of original LDR because segmentation of low-light or extremely bright images does not yield good results. We use Segment Anything (SAM) [50] to generate segmentation classes for the histogram-equalized LDR and reconstructed tone-mapped HDR images (Sec. S2.3 for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as:
$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (5)
where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$. The mean IoU over all segmentation classes can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (6)
We define the semantic segmentation loss as:
$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (7)
This loss is calculated in both forward and backward cycles with the input and output being swapped (Sec. S2.3 for visual depiction of the semantic segmentation loss).
@@
Cycle Consistency Loss. The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each other. Hence, we also incorporate a cycle consistency loss to prevent mutual contradiction of the learned mappings $G_{Y}$ and $G_{X}$. For each image $x_{i}$ from the LDR domain, the cycle of reconstruction i.e., from LDR to HDR and then back to LDR must result back in the original image $x_{i}$. Hence, we can define the forward cycle consistency as: $x_{i}\rightarrow G_{Y}(x_{i})\rightarrow G_{X}(G_{Y}(x_{i}))\approx x_{i}$. Similarly, the backward cycle consistency can be represented as: $y_{j}\rightarrow G_{X}(y_{j})\rightarrow G_{Y}(G_{X}(y_{j}))\approx y_{j}$. The loss can be formulated as:
@@
where $\|\cdot\|_{1}$ is the L1 norm (Sec. S2.3 for visual depiction of the cycle consistency loss).
Identity Loss. For LDR $\leftrightarrow$ HDR translation tasks, we also found that adversarial and cycle consistency formulations alone cannot preserve the color and hue information. This is due to incorrect mapping of color shades from LDR to HDR domains by the generators stemming from the underlying difference in dynamic ranges. Therefore, we force the generators to replicate an identity mapping by providing target domain images. This loss can be expressed as:
$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (11)
Final Loss. The full objective is expressed in Eq. 12, where $\lambda$ scales the relative importance of the cycle consistency and identity loss. $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [117]. $\alpha$ and $\beta$ are weights for the contrastive and semantic segmentation loss, respectively. Both values are set to $2$ (Sec. S3.3 for ablation experiments for $\alpha$ and $\beta$).
$\begin{split}\mathcal{L}_{\text{full}}=\\ \mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\\ \lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\\ \alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}+\mathcal{L}_{\text{llm}}\end{split}$ (12)
4 Experiments
We discuss the experimental setup details in Sec. S2.4. Datasets. For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [16], NTIRE [79], and HDR-Synth & HDR-Real [64] paired datasets. HDRTV has $1235$ samples for training and $117$ for testing. The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples. HDR-Synth & HDR-Real dataset consists of $20537$ samples. Other paired datasets used in our evaluations are DrTMO [23] ($1043$ samples), Kalantari [44] ($89$ samples), HDR-Eye [51] (46 samples), and LDR-HDR Pair [40] ($176$ samples). We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity. We used the pre-defined train/test sets of HDRTV and NTIRE. For HDR-Synth & HDR-Real, we performed a random $80/20$ split for training and testing in all experiments unless specified otherwise. For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR. For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [10] for datasets with only single-exposed LDR images.
Metrics. We use four metrics to report the results. High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [68] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model. Structural Similarity Index Measure (SSIM) [99, 100, 101] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information. Peak Signal-to-Noise Ratio (PSNR) [31] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^1] Unless specified otherwise, for HDR evaluation PSNR is computed on tone-mapped original and reconstructed HDR image pairs.. Learned Perceptual Image Patch Similarity (LPIPS) [115] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge-based contributions. This metric aligns with the perceptual judgement of humans.
Table 2: HDR reconstruction result. Comparison with supervised (gray) and unsupervised, weakly-supervised and self-supervised (black) learning methods trained and evaluated on the paired datasets HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64]. LP: Supervised (S), unsupervised (US), weakly-supervised (WS), and self-supervised (SS).
@@ 4 Experiments
HDRCNN (ACM TOG’17[22] S 20.12 0.724 0.422 60.12
DrTMO (ACM TOG’17[23] S 20.43 0.711 0.412 60.81
ExpandNet (Eurographics’18[69] S 21.44 0.854 0.451 58.12
FHDR (GlobalSIP’19[47] S 20.84 0.841 0.307 62.45
SingleHDR (CVPR’20[64] S 21.01 0.851 0.402 64.21
Two-stage (CVPRW’21[1] S 34.29 0.856 0.287 61.78
HDR-GAN (IEEE TIP’21[77] S 30.11 0.912 0.411 65.01
KUNet (IJCAI’22[96] S 37.21 0.974 0.051 62.14
Ghost-free (ECCV’22[65] S 40.32 0.966 0.043 63.02
ArtHDR-Net (APSIPA’23[5] S 37.12 0.961 0.321 63.43
HistoHDR-Net (ICIP’24[6] S 38.21 0.968 0.301 65.15
UPHDR-GAN (IEEE TCSVT’22[61] US 39.98 0.965 0.044 63.54
PSENet (WACV’23[76] US 27.35 0.856 0.274 62.89
SingleHDR(W) (WACV’23[55] WS 30.79 0.952 0.055 62.95
GlowGAN-ITM (ICCV’23[94] US 30.19 0.901 0.064 60.05
SelfHDR (ICLR’24[116] SS 39.51 0.972 0.037 64.77
LLM-HDR (Ours) SS 40.11 0.979 0.020 68.51
5 Results
5.1 HDR Reconstruction
For an extensive comparison we selected a combination of methods that utilize different learning paradigms including, HDRCNN [22], DrTMO [23], ExpandNet [69], FHDR [47], SingleHDR [64], Two-stage [1], KUNet [96], HistoHDR-Net [6], HDR-GAN [77], ArtHDR-Net [5], Ghost-free HDR [65], SelfHDR [116], UPHDR-GAN [61], GlowGAN-ITM [94], SingleHDR(W) [55], and PSENet [76]. These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised, weakly-supervised, self-supervised, unpaired learning, and knowledge-conditioned.
Paired Datasets. Tab. 2 summarizes the results of evaluation on HDRTV [16], NTIRE [79], HDR-Synth & HDR-Real [64] paired datasets, the most widely used datasets by the state-of-the-art (Sec. S3.1 for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $5$ unsupervised, weakly-supervised, and self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics. It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion with unpaired samples (Sec. 3).
Mixed Datasets. To further demonstrate the strength of our approach we extended the train set used in the previous experiment (paired data) with: 1) The HDR images from DrTMO [23], Kalantari [44], HDR-Eye [51], LDR-HDR Pair [40], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images. There is no overlap between LDR and HDR in this additional data, i.e., it is an unpaired dataset. The final dataset used for training in this experiment has an approximately 87% overlap between LDR and HDR images. We kept the test set identical to the previous experiment, reported in Tab. 2, in order to compare with the state-of-the-art methods based paired data. Given that the only other methods that support this type of unpaired train data are UPHDR-GAN [61] and GlowGAN-ITM [94], we perform a direct comparison with these approaches in Tab. 3. The results demonstrate that our method outperforms all unpaired and paired data methods, setting new state-of-the-art on all metrics (Sec. S3.1 for more mixed data experiments).
Table 3: HDR reconstruction results. Comparison with unsupervised learning methods trained on mixed datasets.
@@ 5.1 HDR Reconstruction
GlowGAN-ITM [94] 33.45 0.910 0.061 62.13
LLM-HDR (Ours) 43.96 0.989 0.020 69.32
Table 4: HDR reconstruction results. Comparison with unsupervised learning methods trained on unpaired datasets.
[…]
UPHDR-GAN [61] 37.51 0.941 0.101 61.63
GlowGAN-ITM [94] 32.15 0.923 0.137 60.11
LLM-HDR (Ours) 39.32 0.952 0.082 64.32
Unpaired Datasets. We performed cross-dataset experiment where LDR and HDR for training are sourced from different datasets. We kept the test set identical to the previous experiments. LDR images are sourced from HDRTV [16] and NTIRE [79], while the HDR are taken from HDR-Synth & HDR-Real [64]. The direct comparison with UPHDR-GAN [61] and GlowGAN-ITM [94] which have unpaired training capabilities is reported in Tab. 4. The results demonstrate that our method outperforms both unpaired data methods on all metrics.
5.2 LDR Reconstruction
We report results for LDR reconstruction in Tab. 5 and compare with the state-of-the-art tone-mapping operators $\mu$-law [43], Reinhard’s [84], and Photomatix [34]. We also consider the HDR-to-LDR prior from GlowGAN [94] which is used to re-project reconstructed HDR to LDR space for inverse learning. The test data is identical to the one used in the HDR reconstruction experiments. The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS. In terms of PSNR, it is second best after Photomatix (Sec. S3.2 for more LDR reconstruction experiments).
Table 5: LDR reconstruction results. Comparison with the state-of-the-art tone-mapping operators and GlowGAN-prior [94].